OpenStudy (anonymous):

how do u factor 9x^2-3x-6 completely ?

OpenStudy (anonymous):

answer = (3x+2)(x-1)

OpenStudy (mathmagician):

at first, find the roots of equation using formula: if you have equation ax^2+bx+c=0, the roots are: x1=(-a+sqrt(b^2-4ac))/2a;x2=(-a-sqrt(b^2-4ac))/2a Then you can factor with formula ax^2+bx+c=a(x-x1)(x-x2)

OpenStudy (anonymous):

=(9x-6) (x+1)

OpenStudy (anonymous):

x=-1 and x=-2/3